# X-ray View of Little Red Dots: Do They Host Supermassive Black Holes?

### v1

@@ Abstract
The discovery of Little Red Dots (LRDs) – a population of compact, high-redshift, dust-reddened galaxies – is one of the most surprising results from JWST. However, the nature of LRDs is still debated: some studies suggest that these galaxies host accreting supermassive black holes (SMBHs), while others conclude that the near-infrared emission primarily originates from intense star formation. In this work, we utilize ultra-deep Chandra observations and study LRDs residing behind the lensing galaxy cluster, Abell 2744. We probe the X-ray emission from individual galaxies but find that they remain undetected and provide SMBH mass upper limits of assuming Eddington limited accretion. To increase the signal-to-noise ratios, we conduct a stacking analysis of the full sample with a total lensed exposure time of Ms. We also bin the galaxies based on their stellar mass, lensing magnification, and detected broad-line H emission. All but one stacked sample remains undetected with SMBH mass upper limits of . We obtain a tentative, detection for LRDs exhibiting broad-line H emission. Taking this detection at face value, the inferred mean SMBH mass is assuming Eddington-limited accretion, about 1.5 orders of magnitude lower than that inferred from JWST data. Our results imply that LRDs do not host over-massive SMBHs and/or accrete at a few percent of their Eddington limit. The significant discrepancy between the JWST and Chandra data hints that the scaling relations used to infer the SMBH mass from the H line and virial relations may not be applicable for high-redshift LRDs.
@@ 1. Introduction
Since its launch, JWST has revolutionized the picture of the early universe on multiple forefronts. In particular, JWST has discovered an unexpectedly large sample of high-redshift galaxies in multiple wide fields and behind lensing galaxy clusters (e.g. Robertson et al., 2023; Castellano et al., 2022a, b; Atek et al., 2023). JWST has also brought into view high-redshift () accreting SMBHs that shed new light on our understanding of early seeding and growth (Bogdán et al., 2024; Natarajan et al., 2024; Kokorev et al., 2023; Kovacs et al., 2024; Maiolino et al., 2024). In addition, thanks to its near-infrared sensitivity, JWST unveiled a previously unknown, dust-reddened, galaxy population at high () redshifts (Labbé et al., 2023; Labbe et al., 2023; Matthee et al., 2023; Akins et al., 2023; Barro et al., 2024). These galaxies, dubbed as Little Red Dots (LRDs), are red, which suggests heavy obscuration, are compact (), and many of them have high stellar masses ().
@@
The discovery of LRDs has triggered a debate about their nature. The emission in LRDs could be either dominated by flux from active galactic nuclei (AGN) or from a population of young stars associated with vigorous star formation. Alternatively, LRDs could be a heterogeneous galaxy population, in which both AGN and star formation contribute to their fluxes to different degrees. To constrain the origin of LRDs, several systems were followed up with JWST, revealing evidence for both of these possibilities. For example, Kocevski et al. (2023) studied the JWST spectrum of two LRDs and found broad H emission, indicating the presence of SMBHs with . Similar spectroscopic follow-up studies were performed for several other LRDs, which also suggest that they could host SMBHs with (Greene et al., 2023; Killi et al., 2023; Matthee et al., 2023). The high SMBH masses imply that the black hole-to-stellar mass ratio of LRDs is much higher than that in the local universe (Magorrian et al., 1998; Häring & Rix, 2004; Kormendy & Ho, 2013; Reines & Volonteri, 2015) or for AGN observed up to (Suh et al., 2020). This result is hardly unexpected: previous Chandra/JWST studies found several high-redshift galaxies that host extremely massive SMBHs with some of them reaching black hole-to-stellar mass ratios of (e.g. Bogdán et al., 2024; Goulding et al., 2023; Übler et al., 2023; Harikane et al., 2023; Kovacs et al., 2024). However, other studies, based on JWST MIRI spectra, suggest that the emission in several LRDs is dominated by young stellar populations. Specifically, modeling the spectral energy distribution of 31 LRDs indicates that their characteristics can be described by intense and compact starbursts and young stellar ages, where the energy output is dominated by the emission from OB stars (Pérez-González et al., 2024).
@@
Although JWST carried out several follow-up observations with the NIRSpec and MIRI instruments to better understand the nature of these galaxies, interpreting the near-infrared data is often not trivial because AGN and star-forming processes could produce very similar spectral signatures. X-ray data, in contrast, can provide important clues about the nature of LRDs for several reasons. First, the X-ray emission originating from dust-obscured AGN emission is much less impacted by obscuration, particularly at high redshifts (Bogdán et al., 2022, 2024; Kovacs et al., 2024). Second, the level of expected X-ray emission from AGN and star-forming processes is significantly different with the latter being several orders of magnitude fainter. Therefore, given the typical exposure times of deep X-ray observations and the sensitivity of present-day X-ray telescopes, emission from star-forming processes will likely remain undetected, while accreting SMBHs could be detected even at large cosmic distances. Finally, the ultra-deep Chandra observations of Abell 2744, combined with the lensing magnification of the galaxy cluster, allow the detection of accreting SMBHs with even at (Bogdán et al., 2024; Kovacs et al., 2024). These SMBH masses are within the range of those predicted by JWST spectroscopic studies of LRDs (Greene et al., 2023; Matthee et al., 2023). Therefore, the Chandra X-ray observations of Abell 2744 are ideally suited to independently probe AGN in LRDs.
@@ 2. Sample of Little Red Dots
We study a sample of high-redshift LRDs in the Abell 2744 field that were detected by JWST (Labbe et al., 2023). From the parent sample of 35 red and compact galaxies, Labbe et al. (2023) identified 26 objects that exhibit red colors (), a compact nature with a median size of pc, and a light profile that is dominated by a central point source-like component. Follow-up spectroscopic studies of these 26 systems demonstrated that three of them are nearby brown dwarfs(Burgasser et al., 2024) and three of them are associated with a triply imaged lensed galaxy at (Furtak et al., 2024). This leaves us with 21 individual LRDs that represent our main sample. Within this sample, Labbe et al. (2023) identified an AGN subset of 12 galaxies whose broadband SED modeling required an AGN component. A sub-sample of LRDs was followed up by JWST spectroscopy (Greene et al., 2023), which indicated that nine LRDs in the Abell 2744 field host AGN with masses of .
The redshifts of the 21 individual LRDs are in the range of with a median of . Of the 21 galaxies, 11 have precise spectroscopic redshifts (Greene et al., 2023; Burgasser et al., 2024; Furtak et al., 2024). The stellar masses of the galaxies span three orders of magnitude and are in the range of with a median of . Thanks to the lensing magnification of Abell 2744, the light from these galaxies is amplified by factors of . The main properties of the galaxies are listed in Table LABEL:tab:sample and for details we refer to previous works (Labbe et al., 2023; Greene et al., 2023).
@@ 4.1. Individual galaxies
The LRDs in our sample were identified as potential AGN host galaxies (Labbe et al., 2023; Greene et al., 2023). Since SMBHs with masses of accreting at their Eddington limit could be detectable in the Abell 2744 field with the deep Chandra observations, we first carry out X-ray photometry at the position of individual LRDs.
@@ 5. Discussion
The main finding of our analysis is the non-detection of AGN in individual LRDs and the weak/non-detection of stacked galaxy samples in the Chandra X-ray data. Given the deep Chandra observations, the lensing magnification of galaxies, and the stacking approach, we obtain upper limits of the order of for typical SMBH masses (Tables LABEL:tab:sample and LABEL:tab:limits). Using these highly constraining limits, we place the population of LRDs on the relation. In Figure 3, we present this relation for the individual and the stacked sample of LRDs. In the plot, we only highlight the limit for all stacked galaxies but the limits on other stacked sub-samples are comparable (Table LABEL:tab:limits). We note that the X-ray upper limits do not show a dependence on the stellar mass of the galaxies, since the limits are primarily driven by the number of counts in the source region, which, in turn, is determined by the depth of the Chandra exposure time and the brightness of the ICM of Abell 2744 at that spatial location.
@@
The discovery of SMBHs in the high-redshift () universe suggests that some of these SMBHs likely originated from heavy seeds, i.e. from the direct collapse of massive gas clouds (Natarajan et al., 2024). However, LRDs reside at lower redshifts and their mean inferred SMBH mass does not exceed . Therefore, when placing these galaxies on BH growth tracks (e.g. Kovacs et al., 2024), we conclude that SMBHs of LRDs could have originated from light seeds, i.e. from the collapse of the first generation of stars, with initial masses of . Our results support the earlier theoretical suggestion that multiple seeding mechanisms likely operate simultaneously in the early universe and that offsets from the local BH mass stellar mass relation at higher redshifts may, in fact, hold important clues to the nature of early BH seeds Natarajan (2012). Models tracing the assembly history of black holes over cosmic time, in which seeding ceases by typically report that even then some seeding signatures may persist to late times Ricarte & Natarajan (2018). While LRDs appear to be consistent with originating from low mass seeds, however, if seed formation does not cease and continues to later cosmic epochs, as proposed by Natarajan (2021), then they could also have originated from late-forming heavy seeds and we could be witnessing extended seeding. At the highest redshifts , massive seeding is predicted to result in the production of a transient class of galaxies, the over-massive black hole galaxies (OBGs) that are offset in the BH mass - stellar mass scaling relation compared to the local one Agarwal et al. (2013) reflecting initial conditions. At later times, if seeding persists, then where BHs end up on the scaling relations may well be determined primarily once again by seeding rather than their accretion history. The prospects for addressing these important open questions looks bright with more detailed follow-up studies of the extended environment of LRDs.
@@
Acknowledgements. TTA acknowledges support from ADAP grant 80NSSC24K0692. Á.B. acknowledges support from the Smithsonian Institution and the Chandra Project through NASA contract NAS8-03060. O.E.K is supported by the GAČR EXPRO grant No. 21-13491X. P.N. acknowledges support from the Gordon and Betty Moore Foundation and the John Templeton Foundation that fund the Black Hole Initiative (BHI) at Harvard University where she serves as one of the PIs.

### v2

@@ Abstract
The discovery of Little Red Dots (LRDs) – a population of compact, high-redshift, dust-reddened galaxies – is one of the most surprising results from JWST. However, the nature of LRDs is still debated: does the near-infrared emission originate from accreting supermassive black holes (SMBHs), or intense star formation? In this work, we utilize ultra-deep Chandra observations and study LRDs residing behind the lensing galaxy cluster, Abell 2744. We probe the X-ray emission from individual galaxies but find that they remain undetected and provide SMBH mass upper limits of assuming Eddington limited accretion. To increase the signal-to-noise ratios, we conduct a stacking analysis of the full sample with a total lensed exposure time of Ms. We also bin the galaxies based on their stellar mass, lensing magnification, and detected broad-line H emission. For the LRDs exhibiting broad-line H emission, there is a hint of a stacked signal (), corresponding to a SMBH mass of . Assuming unobscured, Eddington-limited accretion, this BH mass is at least 1.5 orders of magnitude lower than that inferred from virial mass estimates using JWST spectra. Given galaxy-dominated stellar mass estimates, our results imply that LRDs do not host over-massive SMBHs and/or accrete at a few percent of their Eddington limit. However, alternative stellar mass estimates may still support that LRDs host over-massive BHs. The significant discrepancy between the JWST and Chandra data hints that the scaling relations used to infer the SMBH mass from the H line and virial relations may not be applicable for high-redshift LRDs.
@@ 1. Introduction
Since its launch, JWST has revolutionized the picture of the early universe on multiple forefronts. In particular, JWST has discovered an unexpectedly large sample of high-redshift galaxies in multiple wide fields and behind lensing galaxy clusters (e.g. Robertson et al., 2023; Castellano et al., 2022a, b; Atek et al., 2023). JWST has also brought into view high-redshift () accreting SMBHs that shed new light on our understanding of early seeding and growth (Bogdán et al., 2024; Natarajan et al., 2024; Kokorev et al., 2023; Kovacs et al., 2024; Maiolino et al., 2024a). In addition, thanks to its near-infrared sensitivity, JWST unveiled a previously unknown, dust-reddened, galaxy population at high () redshifts (Labbé et al., 2023; Labbe et al., 2023; Matthee et al., 2023; Akins et al., 2023a; Barro et al., 2024a). These galaxies, dubbed as Little Red Dots (LRDs), are red, which suggests heavy obscuration, are compact (), and many of them have high inferred stellar masses ().
@@
The discovery of LRDs has triggered a debate about their nature. The emission in LRDs could be either dominated by flux from active galactic nuclei (AGN) or from a population of young stars associated with vigorous star formation. Alternatively, LRDs could be a heterogeneous galaxy population, in which both AGN and star formation contribute to their fluxes to different degrees. To constrain the origin of LRDs, several systems were followed up with JWST, revealing evidence for both of these possibilities. For example, Kocevski et al. (2023) studied the JWST spectrum of two LRDs and found broad H emission, indicating the presence of SMBHs with . Similar spectroscopic follow-up studies were performed for several other LRDs, which also suggest that they could host SMBHs with (Greene et al., 2024; Killi et al., 2023; Matthee et al., 2023). The high SMBH masses imply that the black hole-to-stellar mass ratio of LRDs is much higher than that in the local universe (Magorrian et al., 1998; Häring & Rix, 2004; Kormendy & Ho, 2013; Reines & Volonteri, 2015) or for AGN observed up to (Suh et al., 2020). This result is hardly unexpected: previous Chandra/JWST studies found several high-redshift galaxies that host extremely massive SMBHs with some of them reaching black hole-to-stellar mass ratios of (e.g. Bogdán et al., 2024; Goulding et al., 2023; Übler et al., 2023; Harikane et al., 2023; Kovacs et al., 2024). However, other studies, based on JWST MIRI imaging, suggest that the emission in several LRDs is dominated by young stellar populations. Specifically, modeling the spectral energy distribution of 31 LRDs indicates that their characteristics can be described by intense and compact starbursts and young stellar ages, where the energy output is dominated by the emission from OB stars (Pérez-González et al., 2024). An alternative interpretation using panchormatic MIRI photometry and ALMA data is that these are post-starburst galaxies with moderate dust level (Williams et al., 2023).
@@
Although JWST carried out several follow-up observations with the NIRSpec and MIRI instruments to better understand the nature of these galaxies, interpreting the near-infrared data is often not trivial because AGN and star-forming processes could produce very similar spectral signatures. X-ray data, in contrast, can provide important clues about the nature of LRDs for several reasons. First, the X-ray emission from dust-obscured AGN emission is less affected by obscuration. Although a larger fraction of AGN are obscured at higher redshifts, the shift to higher rest-frame X-ray energies allows the detection of heavily obscured AGN (e.g. Bogdán et al., 2022, 2024; Kovacs et al., 2024). Second, the level of expected X-ray emission from AGN and star-forming processes is significantly different with the latter being several orders of magnitude fainter. Therefore, given the typical exposure times of deep X-ray observations and the sensitivity of present-day X-ray telescopes, emission from star-forming processes will likely remain undetected, while accreting SMBHs could be detected even at large cosmic distances. Finally, the ultra-deep Chandra observations of Abell 2744, combined with the lensing magnification of the galaxy cluster, allow the detection of accreting SMBHs with even at (Bogdán et al., 2024; Kovacs et al., 2024). These SMBH masses are within the range of those predicted by JWST spectroscopic studies of LRDs (Greene et al., 2024; Matthee et al., 2023). Therefore, the Chandra X-ray observations of Abell 2744 are ideally suited to independently probe AGN in LRDs.
@@ 2. Sample of Little Red Dots
We study a sample of high-redshift LRDs in the Abell 2744 field that were detected by JWST (Labbe et al., 2023). From the parent sample of 35 red and compact galaxies, Labbe et al. (2023) identified 26 objects that exhibit red colors (), a compact nature with a median size of pc, and a light profile that is dominated by a central point source-like component. Follow-up spectroscopic studies of these 26 systems demonstrated that three of them are nearby brown dwarfs(Burgasser et al., 2024) and three of them are associated with a triply imaged lensed galaxy at (Furtak et al., 2024). This leaves us with 21 individual LRDs that represent our main sample. Within this sample, Labbe et al. (2023) identified an AGN subset of 12 galaxies whose broadband SED modeling required an AGN component. A sub-sample of LRDs was followed up by JWST spectroscopy (Greene et al., 2024), which indicated that nine LRDs in the Abell 2744 field host AGN with masses of .
The redshifts of the 21 individual LRDs are in the range of with a median of . Of the 21 galaxies, 11 have precise spectroscopic redshifts (Greene et al., 2024; Burgasser et al., 2024; Furtak et al., 2024). The stellar masses of the galaxies span three orders of magnitude and are in the range of with a median of . Thanks to the lensing magnification of Abell 2744, the light from these galaxies is amplified by factors of with a median of . The main properties of the galaxies are listed in Table LABEL:tab:sample and for details we refer to previous works (Labbe et al., 2023; Greene et al., 2024).
@@ 3.2. Analysis with StackFast
3.3. Stellar Mass Estimates
In this work, the stellar mass estimates are taken from Labbe et al. (2023). Stellar mass estimates in Labbe et al. (2023) are calculated using SED fitting, and the reported stellar masses are a result of galaxy-only templates fitted to the data. In § 3.3 of Labbe et al. (2023), it is stated that sources that have ALMA data are better fitted with a dominant AGN component. Several other related works have pointed out Akins et al. (2023b); Barro et al. (2024b); Wang et al. (2024) that including an AGN component can lead to stellar mass being lowered by 1-2 magnitudes. A galaxy+AGN template fit can therefore lower the stellar mass estimates used in this work, and will have implications for black hole mass to stellar mass ratio, which is further discussed in the § 5.
@@ 4.1. Individual galaxies
The LRDs in our sample were identified as potential AGN host galaxies (Labbe et al., 2023; Greene et al., 2024). Since SMBHs with masses of accreting at their Eddington limit could be detectable in the Abell 2744 field with the deep Chandra observations, we first carry out X-ray photometry at the position of individual LRDs.
@@ 5. Discussion
The main finding of our analysis is the non-detection of AGN in individual LRDs and the weak/non-detection of stacked galaxy samples in the Chandra X-ray data. Our results are consistent with the X-ray stacking results from Yue et al. (2024), which also find very weak X-ray emission for a set of 19 LRDs with broad H emission, and conclude that LRDs might have different properties compared to typical low-redshift Type-1 AGN. The same X-ray weakness was also observed by Maiolino et al. (2024b), which studied a sample of 71 JWST-selected AGN and concluded that their X-ray weakness indicates that these AGN are Compton-thick.
Given the deep Chandra observations for our sample, the lensing magnification of galaxies, and the stacking approach, we obtain upper limits of the order of for typical SMBH masses (Tables LABEL:tab:sample and LABEL:tab:limits). Using these highly constraining limits, we place the population of LRDs on the relation. In Figure 3, we present this relation for the individual and the stacked sample of LRDs. In the plot, we only highlight the limit for all stacked galaxies but the limits on other stacked sub-samples are comparable (Table LABEL:tab:limits). We note that the X-ray upper limits do not show a dependence on the stellar mass of the galaxies, since the limits are primarily driven by the number of counts in the source region, which, in turn, is determined by the depth of the Chandra exposure time and the brightness of the ICM of Abell 2744 at that spatial location.
@@
An additional caveat is that the stellar mass measurements used in this work are obtained from Labbe et al. (2023), which only provides masses determined from galaxy-only SED fits. However, Labbe et al. (2023) also notes that including ALMA data prefers galaxy+AGN template fits rather than galaxy-only SED fits. Akins et al. (2023b); Barro et al. (2024b); Wang et al. (2024) find AGN signatures in similar high-redshift red objects, and Kocevski et al. (2024) find that for LRDs, the rest-frame optical emission likely originates from AGN, while the rest-frame ultraviolet emission originates from the host galaxy. Barro et al. (2024b) show that if the AGN component is accounted for in mass measurements, the stellar mass could be orders of magnitude lower. If we assume that AGN-dominated template fits will lower stellar masses by orders of magnitude (illustrated by the dashed leftward arrow in Fig. 3), then our results will be consistent with those of galaxies hosting over-massive black holes.
@@
The discovery of SMBHs in the high-redshift () universe suggests that some of these SMBHs likely originated from heavy seeds, i.e. from the direct collapse of massive gas clouds (Natarajan et al., 2024). However, LRDs reside at lower redshifts and their mean inferred SMBH mass does not exceed . Therefore, when placing these galaxies on BH growth tracks (e.g. Kovacs et al., 2024), we conclude that SMBHs of LRDs could have originated from light seeds, i.e. from the collapse of the first generation of stars, with initial masses of . Our results support the earlier theoretical suggestion that multiple seeding mechanisms likely operate simultaneously in the early universe and that offsets from the local BH mass stellar mass relation at higher redshifts may hold important clues to the nature of early BH seeds Natarajan (2012). Models tracing the assembly history of black holes over cosmic time, in which seeding ceases by typically report that even then some seeding signatures may persist to late times Ricarte & Natarajan (2018). While LRDs appear to be consistent with originating from low-mass seeds if seed formation does not cease and continues to later cosmic epochs, as proposed by Natarajan (2021), then they could also have originated from late-forming heavy seeds and we could be witnessing extended seeding. At the highest redshifts , massive seeding is predicted to result in the production of a transient class of galaxies, the over-massive black hole galaxies (OBGs) that are offset in the BH mass - stellar mass scaling relation compared to the local one Agarwal et al. (2013) reflecting initial conditions. At later times, if seeding persists, then where BHs end up on the scaling relations may well be determined primarily once again by seeding rather than their accretion history. The prospects for addressing these important open questions look bright with more detailed follow-up studies of the extended environment of LRDs.
@@
Acknowledgements. We thank the referee for their constructive report. We thank Andy Goulding for the helpful discussions. T.T.A. acknowledges support from ADAP grant 80NSSC24K0692. Á.B. acknowledges support from the Smithsonian Institution and the Chandra Project through NASA contract NAS8-03060. O.E.K is supported by the GAČR EXPRO grant No. 21-13491X. P.N. acknowledges support from the Gordon and Betty Moore Foundation and the John Templeton Foundation that fund the Black Hole Initiative (BHI) at Harvard University where she serves as one of the PIs. This paper employs a list of Chandra datasets, obtained by the Chandra X-ray Observatory, contained in the Chandra Data Collection (CDC) 248 doi:10.25574/cdc.248.
Appendix A Appendix A
For our stacking analysis, photons are extracted by taking into account the value for each source-observation pair. However, in the simplified representation in the co-added image shown in Figure 2, the source aperture size () represents the median for all source-observation pairs for our main sample, the distribution of which is shown in Figure 4. Chandra ACIS-I PSF is described in detail in Chandra PSF library.
Appendix B Appendix B
The analyzed set of Chandra ACIS-I observations are listed in LABEL:tab:data.
| Observation ID | Detector | Observation Date | Observation ID | Detector | Observation Date | ||
| 7915 | 18.11 | ACIS-I | 2006-11-08 | 25278 | 9.78 | ACIS-I | 2022-12-02 |
| 8477 | 44.63 | ACIS-I | 2007-06-10 | 27575 | 19.65 | ACIS-I | 2022-12-02 |
| 8557 | 27.30 | ACIS-I | 2007-06-14 | 25936 | 12.92 | ACIS-I | 2023-01-26 |
| 7712 | 8.07 | ACIS-I | 2007-09-10 | 27678 | 12.42 | ACIS-I | 2023-01-27 |
| 26280 | 11.39 | ACIS-I | 2022-01-18 | 25939 | 14.32 | ACIS-I | 2023-01-28 |
| 25912 | 15.10 | ACIS-I | 2022-04-18 | 27679 | 11.93 | ACIS-I | 2023-01-28 |
| 25911 | 16.59 | ACIS-I | 2022-04-19 | 27680 | 13.21 | ACIS-I | 2023-01-28 |
| 25934 | 18.96 | ACIS-I | 2022-04-21 | 27681 | 9.78 | ACIS-I | 2023-01-29 |
| 25931 | 14.55 | ACIS-I | 2022-04-23 | 25909 | 19.33 | ACIS-I | 2023-05-24 |
| 25954 | 13.39 | ACIS-I | 2022-04-24 | 27856 | 15.88 | ACIS-I | 2023-05-25 |
| 25928 | 15.36 | ACIS-I | 2022-05-03 | 27857 | 12.92 | ACIS-I | 2023-05-26 |
| 25942 | 15.18 | ACIS-I | 2022-05-04 | 27563 | 11.70 | ACIS-I | 2023-06-08 |
| 25958 | 11.81 | ACIS-I | 2022-05-04 | 25941 | 32.65 | ACIS-I | 2023-06-09 |
| 25971 | 12.62 | ACIS-I | 2022-05-04 | 27896 | 13.73 | ACIS-I | 2023-06-10 |
| 25932 | 14.08 | ACIS-I | 2022-05-05 | 25917 | 35.62 | ACIS-I | 2023-06-22 |
| 25972 | 31.14 | ACIS-I | 2022-05-18 | 25950 | 29.69 | ACIS-I | 2023-06-30 |
| 25970 | 23.99 | ACIS-I | 2022-06-12 | 25946 | 29.69 | ACIS-I | 2023-07-01 |
| 25919 | 25.03 | ACIS-I | 2022-06-13 | 25965 | 35.63 | ACIS-I | 2023-07-07 |
| 25920 | 29.67 | ACIS-I | 2022-06-13 | 25960 | 24.76 | ACIS-I | 2023-07-08 |
| 25922 | 31.11 | ACIS-I | 2022-06-14 | 25926 | 61.18 | ACIS-I | 2023-07-12 |
| 25968 | 26.92 | ACIS-I | 2022-07-12 | 25955 | 43.42 | ACIS-I | 2023-07-20 |
| 25967 | 33.12 | ACIS-I | 2022-08-01 | 25921 | 16.87 | ACIS-I | 2023-08-04 |
| 25929 | 26.42 | ACIS-I | 2022-08-26 | 25959 | 15.39 | ACIS-I | 2023-08-05 |
| 25925 | 23.41 | ACIS-I | 2022-09-02 | 27974 | 28.71 | ACIS-I | 2023-08-05 |
| 25956 | 13.90 | ACIS-I | 2022-09-02 | 25940 | 27.72 | ACIS-I | 2023-08-10 |
| 25913 | 19.64 | ACIS-I | 2022-09-03 | 25966 | 18.84 | ACIS-I | 2023-08-13 |
| 25915 | 20.31 | ACIS-I | 2022-09-03 | 28370 | 20.73 | ACIS-I | 2023-08-13 |
| 25923 | 10.64 | ACIS-I | 2022-09-04 | 25933 | 23.88 | ACIS-I | 2023-08-15 |
| 25279 | 23.69 | ACIS-I | 2022-09-06 | 28483 | 20.22 | ACIS-I | 2023-08-19 |
| 25924 | 21.54 | ACIS-I | 2022-09-07 | 25935 | 24.08 | ACIS-I | 2023-08-20 |
| 25944 | 20.85 | ACIS-I | 2022-09-08 | 27780 | 14.90 | ACIS-I | 2023-08-21 |
| 25957 | 21.29 | ACIS-I | 2022-09-08 | 25943 | 16.69 | ACIS-I | 2023-08-31 |
| 27347 | 21.45 | ACIS-I | 2022-09-09 | 28872 | 13.09 | ACIS-I | 2023-09-01 |
| 25918 | 20.38 | ACIS-I | 2022-09-13 | 25916 | 22.20 | ACIS-I | 2023-09-03 |
| 25953 | 24.24 | ACIS-I | 2022-09-17 | 25964 | 20.32 | ACIS-I | 2023-09-05 |
| 25908 | 21.83 | ACIS-I | 2022-09-23 | 25961 | 18.84 | ACIS-I | 2023-09-09 |
| 27449 | 9.78 | ACIS-I | 2022-09-24 | 28886 | 9.96 | ACIS-I | 2023-09-10 |
| 25910 | 19.31 | ACIS-I | 2022-09-25 | 28887 | 19.85 | ACIS-I | 2023-09-10 |
| 27450 | 9.78 | ACIS-I | 2022-09-26 | 25962 | 21.81 | ACIS-I | 2023-09-11 |
| 25945 | 16.77 | ACIS-I | 2022-09-27 | 25927 | 20.53 | ACIS-I | 2023-09-16 |
| 25948 | 26.80 | ACIS-I | 2022-09-30 | 25947 | 14.90 | ACIS-I | 2023-09-24 |
| 25969 | 26.92 | ACIS-I | 2022-10-09 | 28920 | 15.28 | ACIS-I | 2023-09-25 |
| 25914 | 27.77 | ACIS-I | 2022-10-15 | 25952 | 10.84 | ACIS-I | 2023-09-27 |
| 25907 | 36.80 | ACIS-I | 2022-11-08 | 28934 | 19.83 | ACIS-I | 2023-09-29 |
| 25973 | 18.15 | ACIS-I | 2022-11-11 | 27739 | 21.31 | ACIS-I | 2023-10-01 |
| 25930 | 19.16 | ACIS-I | 2022-11-15 | 25277 | 18.68 | ACIS-I | 2023-10-02 |
| 27556 | 24.64 | ACIS-I | 2022-11-15 | 28951 | 12.90 | ACIS-I | 2023-10-05 |
| 25951 | 28.45 | ACIS-I | 2022-11-18 | 28952 | 13.79 | ACIS-I | 2023-10-08 |
| 25938 | 18.12 | ACIS-I | 2022-11-26 | 28910 | 25.75 | ACIS-I | 2023-10-25 |
| 25963 | 37.08 | ACIS-I | 2022-11-26 | 25949 | 20.82 | ACIS-I | 2023-10-27 |
| 25937 | 29.73 | ACIS-I | 2022-11-27 |